OpenStudy (anonymous):

(1 over 4) + (4x over 5) = (11 over 20)

OpenStudy (anonymous):

do you want to solve this equation? \[\frac{ 1 }{ 4 }+\frac{ 4x }{ 5 }=\frac{ 11 }{ 20 }\]

OpenStudy (anonymous):

yes please..

OpenStudy (anonymous):

Here you go